OpenStudy (anonymous):

i need help finding the domain of this function: r(p)=√2-5p

OpenStudy (anonymous):

did you mean \[\sqrt{(2-5p)}\]

OpenStudy (anonymous):

\[\sqrt{2} -5p\]

OpenStudy (anonymous):

top one

OpenStudy (anonymous):

solve this \[2-5p \ge 0\] because the value under the radical cant be negative

OpenStudy (anonymous):

is it p>2/5 ?

OpenStudy (anonymous):

-2 from both sides \[-5p \ge -2\]

OpenStudy (anonymous):

divide by -5 \[p \le \frac{ 2 }{ 5 }\]

OpenStudy (anonymous):

ok but why is this solved as an inequality?

OpenStudy (anonymous):

because for radical equations the value under the radical sign has to be 0 or greater (because square root of negatives are not real numbers)

OpenStudy (anonymous):

so the domain would be 2/5 ?

OpenStudy (anonymous):

no its expressed as an inequality

OpenStudy (anonymous):

okay so pretend the problem was r(p)=2-5p what would the domain be

OpenStudy (anonymous):

so its \[p \le \frac{ 2 }{ 5 }\]

OpenStudy (anonymous):

if it was that it would be all real numbers

OpenStudy (anonymous):

rp = 2-5p I mean would be all reals for domain

OpenStudy (anonymous):

the reason its an inequality is because radicals have to have number zero or over in the side because radicals of negatives are not real numbers

OpenStudy (anonymous):

oh okay thank you i understand now!

OpenStudy (anonymous):

great